OpenStudy (anonymous):

please help! Domain and range of: a. f(x)= 1/x+3 b. g(x)= sqrt(x+6)

OpenStudy (anonymous):

also h(x)= x^3+2x+5

rishavraj (rishavraj):

is it \[\frac{ 1 }{ x + 3 }~~~~~or ~~~~~~\frac{ 1 }{ x } + 3\]

rishavraj (rishavraj):

if its \[\frac{ 1 }{ x + 3 }\] then the denominator can't be zero. i.e.\[x + 3 \neq 0\] so \[x \neq -3\]

OpenStudy (anonymous):

ok thanks! what about range? and the other ones

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

did u get the domain ????

OpenStudy (anonymous):

for which one?

OpenStudy (anonymous):

i got the first one! @rishavraj

OpenStudy (anonymous):

and i think i got sqrt(x+3) but i'm not sure

rishavraj (rishavraj):

the functn in sqrt can't be negative.... so \[x + 3 \ge 0\]

OpenStudy (anonymous):

i got x≥-3

OpenStudy (anonymous):

oops the question was actually sqrt (x+6) so x≥-6 ?

OpenStudy (anonymous):

or is it x>6? because couldn't it be sqrt(6-6)

rishavraj (rishavraj):

its \[x \ge -6\]

OpenStudy (anonymous):

oh oops i meant -6, but still why isn't it x>-6

OpenStudy (anonymous):

you looking for the range of \[f(x)=\frac{1}{x+3}\]?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

i got all real numbers but not sure

OpenStudy (anonymous):

a fraction is only zero if the numerator is zero

OpenStudy (anonymous):

therefore \[\frac{1}{x+3}\] cannot be zero because the numerator is 1

OpenStudy (anonymous):

range is all real numbers except zero

OpenStudy (anonymous):

oh! oops. thanks! what about sqrt(x+6) and h(x)=x^3+2x+5

OpenStudy (anonymous):

anyone?

rishavraj (rishavraj):

its clear the its domain and range is ALL REAL Numbers... i.e R

OpenStudy (anonymous):

which one are we talking about :/

OpenStudy (anonymous):

\[\sqrt{x+6}\] the square root of anything is never less that zero range \[y\geq 0\]

rishavraj (rishavraj):

about h(x) = x^3 + 2x + 5

OpenStudy (anonymous):

cube goes from \(-\infty\) to \(+\infty\)

OpenStudy (anonymous):

oh, ok thanks! and im still confused about the sqrt one?

OpenStudy (anonymous):

\[y\geq 0\]

OpenStudy (anonymous):

why isn't it y>0?

rishavraj (rishavraj):

bcoz it can be even zero thts why.....

OpenStudy (anonymous):

oh. why is the domain x≥-6

OpenStudy (anonymous):

i thought it could be sqrt0 or no

rishavraj (rishavraj):

see the function in sqrt can never be negative...... so \[x + 6 \geq 0\] \[x \geq -6\]

OpenStudy (anonymous):

oh okay thank you so much!